OpenStudy (kaimartin):

find the inverse of (x)=2^3√(x+2)

OpenStudy (prathamesh_m):

To find the inverse, consider: \[y = 2^{3\sqrt{x+2}}\] now find x in terms of y to get the inverse.

OpenStudy (zzr0ck3r):

Is that a square root? If so, then the function is not one-to-one and thus does not have an inverse.

OpenStudy (samigupta8):

To start with, do you know the way of finding the inverse for a function?

OpenStudy (sachintha):

You can use logarithm to simplify. \(\qquad\log_2y = 3\sqrt{x+2}\) Solve for \(x\) and swap the variables.

OpenStudy (zzr0ck3r):

if you solve for x there, you will not get a function

OpenStudy (zzr0ck3r):

I think it is maybe \(2^{\sqrt[3]{x+2}}\)